# DreamControl: Human-Inspired Whole-Body Humanoid Control for Scene Interaction via Guided Diffusion

### v1

@@ ACKNOWLEDGMENT
We are thankful for discussions and help with experiments from B. Rishi, Y. Patel, D. Narayanan, S. Deolasee.
@@ VII Open and closed states for each task
Left Hand Config
@@
Right Hand Config
@@
Pick ACL — ACL AOR — ACR
Precise Punch ACL — ACL ACR — ACR
Precise Kick ACL — ACL ACR — ACR
Press Button ACL — ACL PBR — PBR
Jump ACL — ACL ACR — ACR
Sit ACL — ACL ACR — ACR
Bimanual Pick ACL — ACL ACR — ACR
@@
AOL — ACL ACR — ACR
Pick from. Ground
@@
ACL — ACL ACR — ACR
Pick and Place ACL — ACL AOR — ACR
Open Drawer ACL — ACL DOR — ACR
Open Door ACL — ACL DOR — ACR
VIII Reference trajectory generation
@@
VIII-A Task-specific prompts
@@
VIII-A1 Pick
For each reference trajectory:- We sample a target point, $p^{\text{target}}=\{p^{x}\in\mathcal{U}(1.0,1.2),p^{y}\in\mathcal{U}(-0.4,0.4),p^{z}=1.1\}$. Target time-step is chosen to be, $t_{p}^{\prime}=50+\lfloor 50(p^{x}-1.0)\rfloor$. Then, we set spatial control signal for wrist as follows:-
$\lambda^{\text{right\_wrist}}_{i}=p^{\text{target}}\forall i\in\{t_{p}^{\prime},...,t_{p}^{\prime}+20\}$
@@ VIII-A1 Pick
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \forall i\in\{t_{p}^{\prime}+20,...,t_{p}^{\prime}+40\}$
@@
$\forall i\in\{t_{p}^{\prime},...,t_{p}^{\prime}+20\}$
VIII-A2 Punch
For each reference trajectory:- We sample a target point, $p^{\text{target}}=\{p^{x}\in\mathcal{U}(1.2,1.5),p^{y}\in\mathcal{U}(-0.2,0.0),p^{z}=\mathcal{U}(1.0,1.5)\}$. Target time-step is chosen to be, $t_{p}^{\prime}=30$. Then, we set spatial control signal for wrist as follows:-
$\lambda^{\text{right\_wrist}}_{i}=p^{\text{target}}\forall i\in\{t_{p}^{\prime}-10,...,t_{p}^{\prime}+10\}$
VIII-A3 Kick
For each reference trajectory:- We sample a target point, $p^{\text{target}}=\{p^{x}\in\mathcal{U}(1.0,1.2),p^{y}=0.0,p^{z}=\mathcal{U}(0.5,1.0)\}$. Target time-step is chosen to be, $t_{p}^{\prime}=30$. Then, we set spatial control signal for right foot as follows:-
$\lambda^{\text{right\_foot}}_{i}=p^{\text{target}}\forall i\in\{t_{p}^{\prime}-13,...,t_{p}^{\prime}+10\}$
VIII-A4 Button Press
For each reference trajectory:- We sample a target point, $p^{\text{target}}=\{p^{x}\in\mathcal{U}(1.4,1.8),p^{y}=\mathcal{U}(-0.4,0.4),p^{z}=\mathcal{U}(1.1,1.2)\}$. Target time-step is chosen to be, $t_{p}^{\prime}=70+\lfloor 50(p^{x}-1.4)\rfloor$. Then, we set spatial control signal for right foot as follows:-
$\lambda^{\text{right\_wrist}}_{i}=p^{\text{target}}\forall i\in\{t_{p}^{\prime},...,t_{p}^{\prime}+20\}$
$\lambda^{\text{right\_wrist}}_{i}={p^{\text{target,x}}-0.2,p^{\text{target,y}},p^{\text{target,z}}}\text{ for }i=t_{p}^{\prime}+40$
VIII-A5 Jump
We sample a trajectory with target point, $p^{\text{target}}=\{p^{x}=1.0,p^{y}=0.0,p^{z}=1.9\}$. Target time-step is chosen to be $t_{p}^{\prime}=50$. We set the spatial control signal for pelvis as: $\lambda^{\text{pelvis}}_{i}=p^{\text{target}}\text{ for }i=t_{p}^{\prime}$
VIII-A6 Sit
We sample a trajectory with target point, $p^{\text{target}}=\{p^{x}=1.3,p^{y}=0.8,p^{z}=0.58\}$. Target time-step is chosen to be $t_{p}^{\prime}=100$. We set the spatial control signal for pelvis as: $\lambda^{\text{pelvis}}_{i}=p^{\text{target}}\text{ for }i=t_{p}^{\prime}$
VIII-A7 Bimanual Pick
We use target point, $p^{\text{target}}=\{p^{x}=0.7,p^{y}=0.0,p^{z}=0.65\}$. Target time-step is chosen to be $t_{p}^{\prime}=98$. We set the spatial control signal for left wrist as follows:-
@@ VIII-A7 Bimanual Pick
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -\frac{0.25i}{t_{p}^{\prime}}\}\forall i\in\{0,...,t_{p}^{\prime}\}$
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{0.25i}{t_{p}^{\prime}}\}\forall i\in\{t_{p}^{\prime},...,t_{p}^{\prime}+98\}$
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -\frac{0.25i}{t_{p}^{\prime}}\}\forall i\in\{0,...,t_{p}^{\prime}\}$
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{0.25i}{t_{p}^{\prime}}\}\forall i\in\{t_{p}^{\prime},...,t_{p}^{\prime}+98\}$
VIII-A8 Pick from ground (side grasp)
We use target point, $p^{\text{target}}=\{p^{x}=0.5,p^{y}=0.5,p^{z}=0.1\}$. Target time-step is $t_{p}^{\prime}=98$. We set the spatial control signal for the left wrist as follows:-
$\lambda^{\text{left\_wrist}}_{i}=\{p^{\text{target,x}},p^{\text{target,y}},p^{\text{target,z}}+0.5-\frac{0.5i}{t_{p}^{\prime}}\}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \forall i\in\{0,...,t_{p}^{\prime}\}$
$\lambda^{\text{left\_wrist}}_{i}=\{p^{\text{target,x}},p^{\text{target,y}},p^{\text{target,z}}-0.5+\frac{0.5i}{t_{p}^{\prime}}\}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \forall i\in\{t_{p}^{\prime},...,t_{p}^{\prime}+98\}$
VIII-A9 Pick from ground (top grasp)
We use target point, $p^{\text{target}}=\{p^{x}=1.0,p^{y}=0.0,p^{z}=0.2\}$. Target time-step is $t_{p}^{\prime}=50$. We set the spatial control signal for the right wrist as follows:-
$\lambda^{\text{right\_wrist}}_{i}=\{p^{\text{target,x}},p^{\text{target,y}},p^{\text{target,z}}\}\text{ for }i=t_{p}^{\prime}$
VIII-A10 Pick and place
We use 2 target points, $p^{\text{target,1}}=\{p^{x}=1.2,p^{y}=-0.15,p^{z}=0.6\}$ and $p^{\text{target,2}}=\{p^{x}=1.2,p^{y}=-0.6,p^{z}=0.6\}$. Target time-steps are $t_{p1}^{\prime}=80$ and $t_{p2}^{\prime}=160$. We set the spatial control signal for the right wrist as follows:-
$\lambda^{\text{right\_wrist}}_{i}=p^{\text{target,1}}\text{ for }i=t_{p1}^{\prime}$
$\lambda^{\text{right\_wrist}}_{i}=p^{\text{target,2}}\text{ for }i=t_{p2}^{\prime}$
VIII-A11 Drawer opening
For drawer opening, the number of trajectories in HumanML dataset are very limited (only $18$ with both “drawer” and “open” keywords in text description), due to which the generated motion was very off when prompted with drawer opening and only giving sparse spatial control signal for wrist at the drawer. To address this, we prompt DreamControl to generate $2$ trajectories and no spatial control signal, one with prompt “stand still” and one with prompt “squat and stay in squat position”. We use the generated motion as initialization and solve for the following optimization problem with gradient descent from the current state. The target points are sampled as $p^{\text{target}}=\{p^{x}\in\mathcal{U}(0.3,0.35),p^{y}\in\mathcal{U}(-0.2,-0.1),p^{z}\in\mathcal{U}(0.4,0.8)\}$ and $t_{p}^{\prime}=40$ if $p^{z}\geq 0.7$ and $t_{p}^{\prime}=50$ if $p^{z}<0.7$.
@@ VIII-B Trajectory filtering
The specific thresholds, $\beta^{\text{torso}}$ and $\beta^{\text{pelvis}}$ for tasks that we do filtering and number of filtered trajectories from total sampled are given in Table X. For all other tasks except open drawer and open door, only $1$ trajectory is manually selected for training which can be easily scaled up with careful engineering for their automated filtering. For open drawer and door tasks, as trajectories are obtained by optimization of right arm, it does not need filtering. It must be noted that these heuristic-based filtering is only required because diffusion-generated trajectories are unfit for tracking. This maybe mainly because diffusion model sees out-of-distribution samples during annealing. With more data, the need for filtering out samples maybe eliminated.
TABLE X: # of filtered trajectories and constants used for filtering
@@ VIII-C Trajectory refinement
Further, we refine motions by disabling movement of the non-functional arm based on the task to avoid it’s unnecessary movement in the reference trajectory. Specifically, we set the joint angles of the left arm and right arm group joints denoted as $G^{\text{left arm}}$ and $G^{\text{right arm}}$ as in Table LABEL: to their default values along all time steps. The default joint angles. The default values are given in Table LABEL:. The specific group that was refined to set to default joint angles for each task is given in Table LABEL:. The tasks not included in table do not have this refinement applied.
VIII-C1 Special case for Pick task
@@
VIII-D Sim2Real
For sim2real trajectory generation and refinement, we follow a similar design to their corresponding sim tasks with some changes in prompt and trajectory refinements that we will discuss in this section
IX RL Training
IX-A Model architecture

### v2

@@ ACKNOWLEDGMENT
We are thankful for discussions and help with experiments from B. Rishi, Y. Patel, D. Narayanan, S. Deolasee, V. Rajesh, G. Moffatt.
@@ VII Open and closed states for each task
Left Hand
Config
@@
Right Hand
Config
@@
Group
set to
default $q$
Pick ACL — ACL AOR — ACR $G^{\text{left arm}}$
Precise Punch ACL — ACL ACR — ACR -
Precise Kick ACL — ACL ACR — ACR -
Press Button ACL — ACL PBR — PBR -
Jump ACL — ACL ACR — ACR -
Sit ACL — ACL ACR — ACR -
Bimanual Pick ACL — ACL ACR — ACR -
@@
AOL — ACL ACR — ACR $G^{\text{right arm}}$
Pick from Ground
@@
ACL — ACL ACR — ACR $G^{\text{left arm}}$
Pick and Place ACL — ACL AOR — ACR -
Open Drawer ACL — ACL DOR — ACR -
Open Door ACL — ACL DOR — ACR -
VIII Keyword-based filtering for FID
We use the following keywords to filter out trajectories that are used for FID calculation to evaluate human-ness of keypoint trajectories followed by Ours against TaskOnly+:-
  • Pick: All motions whose prompts contain “pick” keyword in them
  • Press Button: All motions whose prompts contain “press” keyword and “button” keyword in them
  • Precise Punch: All motions whose prompts contain “punch” keyword in them
  • Precise Kick: All motions whose prompts contain “kick” keyword in them
  • Jump: All motions whose prompts contain “jump” keyword in them
IX Reference trajectory generation
@@
IX-A Task-specific prompts
@@
TABLE X: Text prompts for each task deployed on real robot
Task Prompts
Pick “a person stands in place, grabs the cup from side and lifts up”
Press Button “a person stands in place, presses elevator button”
Bimanual Pick “a person raises the toolbox with both hands”
Squat “a person squats in place and stands up”
IX-A1 Pick
For each reference trajectory:- We sample a target point, $p^{\text{target}}=\{p^{x}\in\mathcal{U}(1.0,1.2),p^{y}\in\mathcal{U}(-0.4,0.4),p^{z}=1.1\}$. Target time-step is chosen to be, $t_{g}^{\prime}=50+\lfloor 50(p^{x}-1.0)\rfloor$. Then, we set spatial control signal for wrist as follows:-
$\lambda^{\text{right\_wrist}}_{i}=p^{\text{target}}\forall i\in\{t_{g}^{\prime},...,t_{g}^{\prime}+20\}$
@@ IX-A1 Pick
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \forall i\in\{t_{g}^{\prime}+20,...,t_{g}^{\prime}+40\}$
@@
$\forall i\in\{t_{g}^{\prime},...,t_{g}^{\prime}+20\}$
IX-A2 Precise Punch
For each reference trajectory:- We sample a target point, $p^{\text{target}}=\{p^{x}\in\mathcal{U}(1.2,1.5),p^{y}\in\mathcal{U}(-0.2,0.0),p^{z}=\mathcal{U}(1.0,1.5)\}$. Target time-step is chosen to be, $t_{g}^{\prime}=30$. Then, we set spatial control signal for wrist as follows:-
$\lambda^{\text{right\_wrist}}_{i}=p^{\text{target}}\forall i\in\{t_{g}^{\prime}-10,...,t_{g}^{\prime}+10\}$
IX-A3 Precise Kick
For each reference trajectory:- We sample a target point, $p^{\text{target}}=\{p^{x}\in\mathcal{U}(1.0,1.2),p^{y}=0.0,p^{z}=\mathcal{U}(0.5,1.0)\}$. Target time-step is chosen to be, $t_{g}^{\prime}=30$. Then, we set spatial control signal for right foot as follows:-
$\lambda^{\text{right\_foot}}_{i}=p^{\text{target}}\forall i\in\{t_{g}^{\prime}-13,...,t_{g}^{\prime}+10\}$
IX-A4 Press Button
For each reference trajectory:- We sample a target point, $p^{\text{target}}=\{p^{x}\in\mathcal{U}(1.4,1.8),p^{y}=\mathcal{U}(-0.4,0.4),p^{z}=\mathcal{U}(1.1,1.2)\}$. Target time-step is chosen to be, $t_{g}^{\prime}=70+\lfloor 50(p^{x}-1.4)\rfloor$. Then, we set spatial control signal for right foot as follows:-
$\lambda^{\text{right\_wrist}}_{i}=p^{\text{target}}\forall i\in\{t_{g}^{\prime},...,t_{g}^{\prime}+20\}$
$\lambda^{\text{right\_wrist}}_{i}={p^{\text{target,x}}-0.2,p^{\text{target,y}},p^{\text{target,z}}}\text{ for }i=t_{g}^{\prime}+40$
IX-A5 Jump
We sample a trajectory with target point, $p^{\text{target}}=\{p^{x}=1.0,p^{y}=0.0,p^{z}=1.9\}$. Target time-step is chosen to be $t_{g}^{\prime}=50$. We set the spatial control signal for pelvis as: $\lambda^{\text{pelvis}}_{i}=p^{\text{target}}\text{ for }i=t_{g}^{\prime}$
IX-A6 Sit
We sample a trajectory with target point, $p^{\text{target}}=\{p^{x}=1.3,p^{y}=0.8,p^{z}=0.58\}$. Target time-step is chosen to be $t_{g}^{\prime}=100$. We set the spatial control signal for pelvis as: $\lambda^{\text{pelvis}}_{i}=p^{\text{target}}\text{ for }i=t_{g}^{\prime}$
IX-A7 Bimanual Pick
We use target point, $p^{\text{target}}=\{p^{x}=0.7,p^{y}=0.0,p^{z}=0.65\}$. Target time-step is chosen to be $t_{g}^{\prime}=98$. We set the spatial control signal for left wrist as follows:-
@@ IX-A7 Bimanual Pick
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -\frac{0.25i}{t_{g}^{\prime}}\}\forall i\in\{0,...,t_{g}^{\prime}\}$
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{0.25i}{t_{g}^{\prime}}\}\forall i\in\{t_{g}^{\prime},...,t_{g}^{\prime}+98\}$
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -\frac{0.25i}{t_{g}^{\prime}}\}\forall i\in\{0,...,t_{g}^{\prime}\}$
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{0.25i}{t_{g}^{\prime}}\}\forall i\in\{t_{g}^{\prime},...,t_{g}^{\prime}+98\}$
IX-A8 Pick from Ground (Side Grasp)
We use target point, $p^{\text{target}}=\{p^{x}=0.5,p^{y}=0.5,p^{z}=0.1\}$. Target time-step is $t_{g}^{\prime}=98$. We set the spatial control signal for the left wrist as follows:-
$\lambda^{\text{left\_wrist}}_{i}=\{p^{\text{target,x}},p^{\text{target,y}},p^{\text{target,z}}+0.5-\frac{0.5i}{t_{g}^{\prime}}\}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \forall i\in\{0,...,t_{g}^{\prime}\}$
$\lambda^{\text{left\_wrist}}_{i}=\{p^{\text{target,x}},p^{\text{target,y}},p^{\text{target,z}}-0.5+\frac{0.5i}{t_{g}^{\prime}}\}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \forall i\in\{t_{g}^{\prime},...,t_{g}^{\prime}+98\}$
IX-A9 Pick from Ground (Top Grasp)
We use target point, $p^{\text{target}}=\{p^{x}=1.0,p^{y}=0.0,p^{z}=0.2\}$. Target time-step is $t_{g}^{\prime}=50$. We set the spatial control signal for the right wrist as follows:-
$\lambda^{\text{right\_wrist}}_{i}=\{p^{\text{target,x}},p^{\text{target,y}},p^{\text{target,z}}\}\text{ for }i=t_{g}^{\prime}$
IX-A10 Pick and Place
We use 2 target points, $p^{\text{target,1}}=\{p^{x}=1.2,p^{y}=-0.15,p^{z}=0.6\}$ and $p^{\text{target,2}}=\{p^{x}=1.2,p^{y}=-0.6,p^{z}=0.6\}$. Target time-steps are $t_{g1}^{\prime}=80$ and $t_{g2}^{\prime}=160$. We set the spatial control signal for the right wrist as follows:-
$\lambda^{\text{right\_wrist}}_{i}=p^{\text{target,1}}\text{ for }i=t_{g1}^{\prime}$
$\lambda^{\text{right\_wrist}}_{i}=p^{\text{target,2}}\text{ for }i=t_{g2}^{\prime}$
IX-A11 Open Drawer
For drawer opening, the number of trajectories in HumanML dataset are very limited (only $18$ with both “drawer” and “open” keywords in text description), due to which the generated motion was very off when prompted with drawer opening and only giving sparse spatial control signal for wrist at the drawer. To address this, we prompt DreamControl to generate $2$ trajectories and no spatial control signal, one with prompt “stand still” and one with prompt “squat and stay in squat position”. We use the generated motion as initialization and solve for the following optimization problem with gradient descent from the current state. The target points are sampled as $p^{\text{target}}=\{p^{x}\in\mathcal{U}(0.3,0.35),p^{y}\in\mathcal{U}(-0.2,-0.1),p^{z}\in\mathcal{U}(0.4,0.8)\}$ and $t_{g}^{\prime}=40$ if $p^{z}\geq 0.7$ and $t_{g}^{\prime}=50$ if $p^{z}<0.7$.
@@ IX-B Trajectory filtering
The specific thresholds, $\beta^{\text{torso}}$ and $\beta^{\text{pelvis}}$ for tasks that we do filtering and number of filtered trajectories from total sampled are given in Table XI. For all other tasks except Open Drawer and Open Door, only $1$ trajectory is manually selected for training which can be easily scaled up with careful engineering for their automated filtering. For Open Drawer and door tasks, as trajectories are obtained by optimization of right arm, it does not need filtering. It must be noted that these heuristic-based filtering is only required because diffusion-generated trajectories are unfit for tracking. This maybe mainly because diffusion model sees out-of-distribution samples during annealing. With more data, the need for filtering out samples maybe eliminated.
TABLE XI: # of filtered trajectories and constants used for filtering
@@ IX-C Trajectory refinement
Further, we refine motions by disabling movement of the non-functional arm based on the task to avoid it’s unnecessary movement in the reference trajectory. Specifically, we set the joint angles of the left arm and right arm group joints denoted as $G^{\text{left arm}}$ and $G^{\text{right arm}}$ as in Table VII to their default values along all time steps. The default joint angles. The default values are given in Table VI. The specific group that was refined to set to default joint angles for each task is given in Table VII. The tasks not included in table do not have this refinement applied.
IX-C1 Special case for Pick task
@@
IX-D Sim2Real
For sim2real trajectory generation and refinement, we follow a similar design to their corresponding sim tasks with some changes in prompt and trajectory refinements that we will discuss in this section. There are $5$ tasks that we put on hardware robot. Specific prompts used for them are listed in Table X. In addition to the standard refinements as made in the sim versions of these tasks (except Squat), the specific spatial control points and refinements for each task are as follows:-
IX-D1 Pick
For each reference trajectory:- We sample a target point, $p^{\text{target}}=\{p^{x}\in\mathcal{U}(0.25,0.35),p^{y}\in\mathcal{U}(-0.3,0.0),p^{z}=1.1\}$. Target time-step is chosen to be, $t_{g}^{\prime}=30$. Then, we set spatial control signal for wrist as follows:-
$\lambda^{\text{right\_wrist}}_{i}=p^{\text{target}}\forall i\in\{t_{g}^{\prime},...,t_{g}^{\prime}+20\}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \forall i\in\{t_{g}^{\prime}+20,...,t_{g}^{\prime}+40\}$
$\forall i\in\{t_{g}^{\prime},...,t_{g}^{\prime}+20\}$
Of the generated trajectories, we refine them to avoid collision with the platform, plus we also add an additional cost to bring the trajectories close to their corresponding goal points for which the trajectories were generated.
$-\lVert p_{t}^{\text{right hand}}-p_{t-1}^{\text{right hand}}\rVert_{2})|$
$+M_{t}\lVert p_{t}^{\text{right wrist}}-\lambda_{t}^{\text{right wrist,'}}\rVert$
$s.t.\ \ d(p_{t}^{\text{right hand}})=0$
Where $M_{t}=1$ for $t\in\{0,t_{g}^{\prime},t_{g}^{\prime}+1,...,t_{g}^{\prime}+40\}\text{ and }0\text{ otherwise}$. This is to mask for timesteps where the goal is specified. Idea is to minimally modify the trajectories such that they follow the corresponding goal. This refinement was required to ensure smooth policies. We also set the lower body (groups $G^{\text{left leg}}$ and $G^{\text{right leg}}$) to fixed default joint angles and adjust the root height such that the feet touches the ground. This is to enforce that the motion is stand still.
IX-D2 Precise Punch
For each reference trajectory:- We sample a target point, $p^{\text{target}}=\{p^{x}\in\mathcal{U}(0.6,0.65),p^{y}\in\mathcal{U}(-0.2,0.1),p^{z}=\mathcal{U}(1.15,1.4)\}$. Target time-step is chosen to be, $t_{g}^{\prime}=30$. Then, we set spatial control signal for wrist as follows:-
$\lambda^{\text{right\_wrist}}_{i}=p^{\text{target}}\forall i\in\{t_{g}^{\prime},...,t_{g}^{\prime}+20\}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \forall i\in\{t_{g}^{\prime}+20,...,t_{g}^{\prime}+40\}$
Of the generated trajectories, we refine them to bring the trajectories close to their corresponding goal points for which the trajectories were generated.
$-\lVert p_{t}^{\text{right hand}}-p_{t-1}^{\text{right hand}}\rVert_{2})|$
$+M_{t}\lVert p_{t}^{\text{right wrist}}-\lambda_{t}^{\text{right wrist,'}}\rVert$
$s.t.\ \ d(p_{t}^{\text{right hand}})=0$
Where $M_{t}=1$ for $t\in\{0,t_{g}^{\prime}\}\text{ and }0\text{ otherwise}$. This is to mask for timesteps where the goal is specified. Idea is to minimally modify the trajectories such that they follow the corresponding goal. This refinement was required to ensure smooth policies. We also set the lower body (groups $G^{\text{left leg}}$ and $G^{\text{right leg}}$) to fixed default joint angles and adjust the root height such that the feet touches the ground. This is to enforce that the motion is stand still.
IX-D3 Bimanual Pick
We follow the exact same spatial control points as the sim to generate the trajectory. We additionally refine to dismiss any feet slipping by using IK to adjust $G^{\text{left leg}}$ and $G^{\text{right foot}}$ to bring left and right foot to their positions at $p_{0}^{\text{left foot}}$ and $p_{0}^{\text{right foot}}$. Other than this, we set the rollroll and yawyaw of root to 0 for symmetry.
IX-D4 Open Drawer
We follow the exact same procedure as the open drawer trajectories for sim.
IX-D5 Squat
For each trajectory, we sample a trajectory with target point, $p^{\text{target}}=\{p^{x}=-0.15,p^{y}=0.0,p^{z}=\mathcal{U}(0.4,0.6)\}$. Target time-step is chosen to be $t_{g}^{\prime}=100$. We set the spatial control signal for pelvis as: $\lambda^{\text{pelvis}}_{i}=p^{\text{target}}\text{ for }i=t_{g}^{\prime}$. We additionally refine to dismiss any feet slipping by using IK to adjust $G^{\text{left leg}}$ and $G^{\text{right foot}}$ to bring left and right foot to their positions at $p_{0}^{\text{left foot}}$ and $p_{0}^{\text{right foot}}$. Other than this, we set the rollroll and yawyaw of root to 0 for symmetry.
X RL Training
X-A Model architecture